itsmehjay:

A rectangular pool has a length that is 13 meters longer than its width. Its perimeter is 74 meters. What is the width of this pool?

snowflake0531:

Set up your equations length is 13 plus width what would the equation be like? set length as l, width is w

itsmehjay:

um................... 13l + w = ? e.e

itsmehjay:

haha idek xd

snowflake0531:

e.e length equals 13 plus width put that in letters and symbols._.

itsmehjay:

eh? l = 13 + w?

snowflake0531:

yep and then we have perimeter equals 74, so 2l+2w=74 So now, we these two equations, solve for w

itsmehjay:

2l + 2w = 74 13 + w = l don't I multiply the second equation by -1 since I have 2 posi- wait are we even eliminating??

itsmehjay:

omg- im sooo lost ;c

snowflake0531:

l=13+w 2l+2w=74 2(13+w) + 2w=74

itsmehjay:

?????

snowflake0531:

You substitute l=13+w into the equation of 2l +2w=74

itsmehjay:

Wait.. 26w + 2w = 74??

itsmehjay:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 You substitute l=13+w into the equation of 2l +2w=74 \(\color{#0cbb34}{\text{End of Quote}}\) ohhhh we were substituting??? e.e

snowflake0531:

e.e ye

itsmehjay:

i thought we were eliminating sorryy

snowflake0531:

so what'd you get as the answer xd

itsmehjay:

ahhhh its about to be lateeee

itsmehjay:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 so what'd you get as the answer xd \(\color{#0cbb34}{\text{End of Quote}}\) oh I got a mixed number um...

snowflake0531:

._. how'd you do it

itsmehjay:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @itsmehjay Wait.. 26w + 2w = 74?? \(\color{#0cbb34}{\text{End of Quote}}\) was this correct ^ like the equation after you distribute?

snowflake0531:

sniff

itsmehjay:

acklvklghltyimkl7iluo

snowflake0531:

2(13+w) + 2w=74 a(b+c) = ab+ac distribute the parenthesis again._.

itsmehjay:

okiee >.>

itsmehjay:

ohhh 26w + 2w.. +2w = 74?????

snowflake0531:

yep and solve for w

itsmehjay:

add all them together- 30w = 74 74/30 = ?

itsmehjay:

IDKKKK

itsmehjay:

It's alr late snow ._.

snowflake0531:

26+4w=74 4x=48 x=12